kane's grandfather deposited some money for him when he was born and it eared 12%p.a. interest, compounded quarterly, untill he turned 21. if he recieved a cheque for $25,000, what was the initial deposit?
Answer by lwsshak3(11628)

kane's grandfather deposited some money for him when he was born and it eared 12%p.a. interest, compounded quarterly, untill he turned 21. if he recieved a cheque for $25,000, what was the initial deposit
..
The formula to use here:
A=P(1+r)^n, where,
A=amt earned
P=Initial investment=what we want to find out
r=interest per period=.12/4=.03
n=number of periods=21*4=84
25000=P(1+.03)^84
25000=P(1.03)^84
P=25000/(1.03)^84
P=$2087.44
ans:
The initial deposit 21 years ago=$2087.44